# onlineReferences

### v5

@@ 4   Extensions
This section contains extensions to non-binary, ordered treatments and repeated cross sections. It also includes the case that the treatment path is introduced in potential outcomes in $2\times 2$ DID-IV designs. For more details and proofs, see online Appendix Section A.
@@
With a non-binary, ordered treatment in staggered DID-IV designs, our target parameter is the cohort specific average causal response on the treated (CACRT), a natural generalization of the ACRT. The identifying assumptions are the same as in Subsection 3.3.
@@
Introducing the treatment path in $2\times 2$ DID-IV designs
In Subsection 2.1, we impose Assumption 1 (no carryover assumption) on potential outcomes $Y_{i,t}(d,z)$, as we consider the non-staggered adoption of the treatment. We can extend our results to the case that we introduce the treatment path in potential outcomes with binary treatment. Online Appendix Section A.3 shows that the similar assumptions to Assumptions 1-6 are sufficient for the Wald-DID to capture the LATET-like parameter.
@@ 5.1   Stacked two stage least squares regression
We create the data sets for each $CLATT_{e,e+l}$. Each data set includes the units of time $t=e-1$ and $t=e+l$, who are either in cohort $e\in\{2,\dots,T\}$ or in the set of some unexposed cohorts, $U$ ($e\notin U$). Researchers should carefully specify the choice set $U$ in practice. If a never exposed cohort $E_{i}=\infty$ exists, we can set $U=\{\infty\}$ and estimate $CLATT_{e,e+l}$ for all cohorts ($e\in\{2,\dots,T\}$) and all relative time periods ($l\geq 0$). If no never exposed cohorts exist, we can set $U=\max\{E_{i}\}$, the last exposed cohort, and can estimate $CLATT_{e,e+l}$ for cohort $e\in\{2,\dots,\max\{E_{i}\}-1\}$ and relative time period $l\in\{0,\dots,\max\{E_{i}\}-1-e\}$ because every unit will be exposed to the instrument after time $t=\max\{E_{i}\}$.
@@
$Y_{i,t}=\beta^{e,l}_{0}+\beta^{e,l}_{i,.}\mathbf{1}\{E_{i}=e\}+\beta^{e,l}_{,.t}\mathbf{1}\{T_{i}=e+l\}+\beta^{e,l}_{IV}D_{i,t}+\epsilon^{e,l}_{i,t},$
@@
Here, $E_{N}[\cdot]$ is the sample analog of the conditional expectation. Note that the IV estimator $\widehat{CLATT}_{e,e+l}$ is the Wald-DID estimator for each $CLATT_{e,e+l}$, where the choice set $U$ is an unexposed group and time $t=e-1$ is a pre-exposed period. From the estimation procedure, we call this a stacked two-stage least squares (STS) estimator. [^18] Our STS estimator is related to the DID estimators in staggered DID designs proposed by Callaway and Sant’Anna (2021) and Sun and Abraham (2021). The DID estimator of the treatment and the outcome in our STS estimator corresponds to that of Sun and Abraham (2021), and coincides with that of Callaway and Sant’Anna (2021) for the case when no covariates exist and never treated units are used as a control group. Their proposed methods avoid the issue of TWFE estimators in staggered DID designs, while our STS estimator avoids the issue of TWFEIV estimators in staggered DID-IV designs.
@@
Suppose Assumptions 8-14 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal.
[…]
Proof.
See Appendix. ∎
@@ 5.2   Weighting scheme
In staggered DID-IV designs, we propose the following summary measure $\theta_{sel}(\tilde{e})^{IV}$ that corresponds with $\theta_{sel}(\tilde{e})$:
@@
where $\hat{w}(e,t)$ is the sample analog of each ${w}(e,t)$ and $\widehat{CLATT}_{e,t}$ is the consistent estimator for each $CLATT_{e,t}$, obtained from the STS regression in Subsection 5.1. Each $\hat{w}(e,t)$ is the regular asymptotically linear estimator:
@@
[…]
@@ 5.3   Pre-trends test
Suppose that data are available for period $t=-1$. Then, one can assess the plausibility of Assumption 5 between period $t=-1$ and $t=0$ by testing the following null hypothesis
@@
Similarly, one can assess the plausibility of Assumption 6 between period $t=-1$ and $t=0$ by testing the following null hypothesis
@@
These tests can be generalized to multiple pre-exposed periods settings by using the pre-trends tests recently developed in the context of DID designs (Callaway and Sant’Anna (2021), Sun and Abraham (2021), Borusyak et al. (2024)); that is, one can apply these tools to the first stage and the reduced form ($D_{i,t}$ or $Y_{i,t}$ is outcome, $Z_{i,t}$ is treatment) respectively to confirm the plausibility of Assumption 13 and 14 in pre-exposed periods.
@@ 6.3   Illustrating our estimation method
The result is shown in Table 6. The TWFEIV estimate is $-0.009$ and not significantly different from 00. This indicates that, on the whole, the returns to schooling in the UK is nearly zero. However, this may be a misleading conclusion. We cannot interpret that the TWFEIV estimand captures the properly weighted average of all possible CACRTs if the effect of the school reform on education attainment or log annual earnings is not stable across cohorts (Miyaji (2024)).
@@
We now employ the proposed method to estimate each CACRT. First, we create data sets by cohort $t$ ($t=1947,\dots,1956$). Each data set only contains units of cohort $t$ and cohort $t=1946$ in England and North Ireland. We define North Ireland as an unexposed group $U$ and discard the units aged at $14$ between $1957$ and $1965$ to avoid the bad comparisons. We then run TSLS regression in each data set. The standard error is calculated by using the influence function shown in equation (B) in online Appendix Section LABEL:ApeC.1.
@@ 7   Conclusion
In this paper, we formalize an instrumented difference-in-differences (DID-IV). First, we consider a simple setting with two periods and two groups. In this setting, our DID-IV design mainly comprises a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between the two groups. We show that in $2\times 2$ DID-IV designs, the Wald-DID estimand captures the local average treatment effect on the treated (LATET). After establishing $2\times 2$ DID-IV designs, we compare DID-IV to Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018), and point out the issues inherent in Fuzzy DID designs.
@@
$E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]$
$=$ $E[Y_{1}(D_{1}((0,1)))-Y_{0}(D_{0}((0,1)))|E=1]-E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=0]$
$=$ $E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1]$
+$+\{E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,1)))|E=1]-E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=0]\}$
$=$ $E[Y_{1}(D_{1}((0,1)))-Y_{1}(D_{1}((0,0)))|E=1]$
+$+\{E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=1]-E[Y_{1}(D_{1}((0,0)))-Y_{0}(D_{0}((0,0)))|E=0]\}$
$=$ $E[(D_{1}((0,1))-D_{1}((0,0)))\cdot(Y_{1}(1)-Y_{1}(0))|E=1]$
$=$ $LATET\cdot Pr(CM^{Z}|E=1).$ (5)
@@
$E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]$
$=$ $E[D_{1}((0,1))-D_{0}((0,1))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\hskip 8.53581pt$
$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
+$+\{E[D_{1}((0,0))-D_{0}((0,1))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\}$
$=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1]$
+$+\{E[D_{1}((0,0))-D_{0}((0,0))|E=1]-E[D_{1}((0,0))-D_{0}((0,0))|E=0]\}$
$=$ $Pr(CM^{Z}|E=1).$ (6)
where the second equality follows from the simple manipulation. The third equality follows from Assumption 4 and the forth equality follows from Assumption 3 and Assumption 5.
@@
First, we prove the consistency of our SLS estimator. We assume that $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ are well defined. More precisely, we assume that cohort $e$ is non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}=e\}>0$ and there exist some sets of the unexposed cohort $U\subseteq\{u:e-1<u\leq T\}$ that are non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}\in U\}>0$. This assumption guarantees that each Wald-DID estimator $\widehat{CLATT}_{e,e+l}$ is also well defined.
First we derive the probability limit of $\widehat{CLATT}_{e,e+l}$. By the Law of Large Numbers and continuous mapping theorem, we have
[…]
Next, we show that the right hand side of (7) is equal to $CLATT_{e,e+l}$. For the numerator, we have
$E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]$
$=$ $E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-\sum_{e\in U}E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]Pr(E_{i}=e|E_{i}\in U)$
$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}=e]$ (8)
$-\sum_{e\in U}E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}\in U]Pr(E_{i}=e|E_{i}\in U)$
$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e+l}(D_{i,e+l}^{\infty})|E_{i}=e]+E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}=e]$
$-\sum_{e\in U}E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}\in U]Pr(E_{i}=e|E_{i}\in U)$
$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e+l}(D_{i,e+l}^{\infty})|E_{i}=e]$
$+E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})]-E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})]$
$=$ $E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}]Pr(CM_{e,e+l}|E_{i}=e).$ (9)
The third equality follows from Assumption 9, 10 and 12. The fourth equality follows from Assumption 14 and the fifth follows from Assumption 11. For the denominator, by similar calculations,
$E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]$
$=E[D_{i,e+l}^{e}-D_{i,e+l}^{\infty}|E_{i}=e]+E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]$ (10)
$-\sum_{e\in U}E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]Pr(E_{i}=e|E_{i}\in U)$
$=Pr(CM_{e,e+l}|E_{i}=e),$ (11)
where the final equality follows from Assumption 11 and Assumption 13. Combining (9) with (11), we obtain the desirable result.
@@
Next, we prove that our STS estimator is asymptotically normal, deriving its influence function. We define $\theta_{e,l}$ to be
@@
$=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}\Bigg{(}\frac{\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]}{E[\mathbf{1}\{E_{i}=e\}]}-\frac{\mathbf{1}\{E_{i}=U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}=U])\right]}{E[\mathbf{1}\{E_{i}=U\}]}\Bigg{)},$ (12)
@@
$+\frac{E\big{[}\mathbf{1}\{E_{i}=U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}=U]\right]^{2}\big{]}}{E[\mathbf{1}\{E_{i}=U\}]^{2}}\Bigg{)}.$

### v6

@@ 4  Extensions
This section contains extensions to non-binary, ordered treatments and repeated cross sections. For more details and proofs, see online Appendix Section A.
@@
With a non-binary, ordered treatment in staggered DID-IV designs, our target parameter is the cohort specific average causal response on the treated (CACRT), a natural generalization of the ACRT. The estimand and the associated identifying assumptions are the same as in Subsection 3.3.
@@ 5.1  Stacked two stage least squares regression
We create the data sets for each $CLATT_{e,e+l}$. Each data set includes the units of time t=e1t=e-1 and $t=e+l$, who are either in cohort $e\in\mathcal{E}$ or in the set of some unexposed cohorts, $U$ ($e\notin U$). If we adopt the parallel trends assumption based on a never-exposed cohort (Assumptions 1314), we can set $U=\{\infty\}$. In this case, we can estimate $CLATT_{e,e+l}$ for all $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$. If we adopt the parallel trends assumption based on not-yet-exposed cohorts (Assumptions 1617), we can set $U=\{e^{\prime}\in\mathcal{S}(E_{i}):e^{\prime}>e+l\}$. In this case, we can estimate $CLATT_{e,e+l}$ for all $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$.
@@
$Y_{i,t}=\beta^{e,l}_{0}+\beta^{e,l}_{1}\mathbf{1}\{E_{i}=e\}+\beta^{e,l}_{2}\mathbf{1}\{T_{i}=e+l\}+\beta^{e,l}_{IV}D_{i,t}+\epsilon^{e,l}_{i,t}.$
The first stage regression is:
$D_{i,t}=\delta^{e,l}_{0}+\delta^{e,l}_{1}\mathbf{1}\{E_{i}=e\}+\delta^{e,l}_{2}\mathbf{1}\{T_{i}=e+l\}+\delta^{e,l}_{3}\mathbf{1}\{E_{i}=e\}\mathbf{1}\{T_{i}=e+l\}+\eta^{e,l}_{i,t}.$
@@
Here, $E_{N}[\cdot]$ is the sample analog of the conditional expectation. From the estimation procedure, we call this a stacked two-stage least squares (STS) estimator. [^17] Our STS estimator is related to the DID estimators in staggered DID designs proposed by Callaway and Sant’Anna (2021) and Sun and Abraham (2021). The DID estimator of the treatment and the outcome in our STS estimator corresponds to that of Sun and Abraham (2021), and coincides with that of Callaway and Sant’Anna (2021) for the case when no covariates exist and never treated units are used as a control group. Their proposed methods avoid the issue of TWFE estimators in staggered DID designs, while our STS estimator avoids the issue of TWFEIV estimators in staggered DID-IV designs.
@@
  • (i)
    Suppose Assumptions 815 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal.
    $\sqrt{n}(\widehat{CLATT}_{e,e+l}-CLATT_{e,e+l})\xrightarrow{d}\mathcal{N}(0,V(\psi_{i,e,l})).$
  • (ii)
    Suppose that Assumptions 812 and 1618 hold. Then, the STS estimator $\widehat{CLATT}_{e,e+l}$ is consistent and asymptotically normal.
    $\sqrt{n}(\widehat{CLATT}_{e,e+l}-CLATT_{e,e+l})\xrightarrow{d}\mathcal{N}(0,V(\psi_{i,e,l})).$
@@ 5.2  Weighting scheme
[…]
@@
where $\hat{w}(e,t)$ is the sample analog of each w(e,t){w}(e,t) and $\widehat{CLATT}_{e,t}$ is the consistent estimator for each $CLATT_{e,t}$, obtained from the STS regression in Subsection 5.1. Each $\hat{w}(e,t)$ is the regular asymptotically linear estimator:
@@
$\sqrt{n}(\hat{\theta}_{IV}-\theta_{IV})\xrightarrow{d}\mathcal{N}(0,V(l_{i}^{\theta_{IV}})),$
@@ 5.3  Pre-trends test
Suppose that data are available for period t=1t=-1. Then, one can assess the plausibility of Assumption 5 between period t=1t=-1 and $t=0$ by testing the following null hypothesis
@@
Similarly, one can assess the plausibility of Assumption 6 between period t=1t=-1 and $t=0$ by testing the following null hypothesis
@@
These tests can be generalized to multiple pre-exposed periods settings by using the pre-trends tests recently developed in the context of DID designs (Callaway and Sant’Anna (2021), Sun and Abraham (2021), Borusyak et al. (2024)); that is, one can apply these tools to the first stage and the reduced form ($D_{i,t}$ or $Y_{i,t}$ is outcome, $Z_{i,t}$ is treatment) respectively to confirm the plausibility of Assumptions 1314 or Assumptions 1617 in pre-exposed periods.
@@ 6.3  Illustrating our estimation method
The result is shown in Table 6. The TWFEIV estimate is 0.009-0.009 and not significantly different from 0. This indicates that, on the whole, the returns to schooling in the UK is nearly zero. [^20] Young (2024) revisits Oreopoulos (2006) and shows that the 2SLS estimates in this setting can be numerically unstable due to near collinearity among cohort indicators. Following Young (2024), we conduct stability checks based on variable and observation ordering, and find that our TWFEIV estimate does not suffer from such numerical instability. However, this may be a misleading conclusion. We cannot interpret that the TWFEIV estimand captures the properly weighted average of all possible CACRTs if the effect of the school reform on education attainment or log annual earnings is not stable across cohorts (Miyaji (2024)).
@@
We now employ the proposed method to estimate each CACRT. First, we create data sets by cohort $t$ ($t=1947,\dots,1956$). Each data set only contains units of cohort $t$ and cohort $t=1946$ in England and North Ireland. We define North Ireland as an unexposed group $U$. We then run the STS regression in each data set. The standard error is calculated by using the influence function shown in equation (B) in Online Appendix Section B.
@@ 7  Conclusion
In this paper, we formalize an instrumented difference-in-differences (DID-IV). First, we consider a simple setting with two periods and two groups. In this setting, our DID-IV design mainly comprises a monotonicity assumption, and parallel trends assumptions in the treatment and the outcome between the two groups. We show that in $2\times 2$ DID-IV designs, the Wald-DID estimand captures the local average treatment effect on the treated (LATET). After establishing $2\times 2$ DID-IV designs, we clarify the differences between DID-IV and Fuzzy DID designs considered in de Chaisemartin and D’Haultfœuille (2018), and provide the implications of these differences for treatment adoption behavior, the interpretation of the target parameter, and the use of the Wald-DID estimand.
@@
$E[Y_{i,1}-Y_{i,0}|E_{i}=1]-E[Y_{i,1}-Y_{i,0}|E_{i}=0]$
$=$ $E[Y_{i,1}(D_{i,1}((0,1)))-Y_{i,0}(D_{i,0}((0,1)))|E_{i}=1]$
- $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0]$
$=$ $E[Y_{i,1}(D_{i,1}((0,1)))-Y_{i,1}(D_{i,1}((0,0)))|E_{i}=1]$
$+$ $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,1)))|E_{i}=1]$
- $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0]$
$=$ $E[Y_{i,1}(D_{i,1}((0,1)))-Y_{i,1}(D_{i,1}((0,0)))|E_{i}=1]$
$+$ $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1]$
- $E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0]$
$=$ $E[(D_{i,1}((0,1))-D_{i,1}((0,0)))\cdot(Y_{i,1}(1)-Y_{i,1}(0))|E_{i}=1]$
$=$ $LATET\cdot Pr(CM^{Z}|E_{i}=1).$ (5)
@@
$E[D_{i,1}-D_{i,0}\mid E_{i}=1]-E[D_{i,1}-D_{i,0}\mid E_{i}=0]$
$=$ $E[D_{i,1}((0,1))-D_{i,0}((0,1))\mid E_{i}=1]-E[D_{i,1}((0,0))-D_{i,0}((0,0))\mid E_{i}=0]$
$=$ $E[D_{i,1}((0,1))-D_{i,1}((0,0))\mid E_{i}=1]$
$+\Big\{E[D_{i,1}((0,0))-D_{i,0}((0,1))\mid E_{i}=1]-E[D_{i,1}((0,0))-D_{i,0}((0,0))\mid E_{i}=0]\Big\}$
$=$ $E[D_{i,1}((0,1))-D_{i,1}((0,0))\mid E_{i}=1]$
$+\Big\{E[D_{i,1}((0,0))-D_{i,0}((0,0))\mid E_{i}=1]-E[D_{i,1}((0,0))-D_{i,0}((0,0))\mid E_{i}=0]\Big\}$
$=$ $\Pr(CM^{Z}\mid E_{i}=1).$ (6)
Here, the second equality follows from the simple manipulation. The third equality follows from Assumption 4 and the forth equality follows from Assumption 3 and Assumption 5.
@@
Fix $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$. Note that Assumption 15 guarantees that the Wald-DID estimand $w_{e,l}^{DID}$ is well defined.
First, we consider the numerator of $w_{e,l}^{DID}$. From Assumptions 815, we have:
$E[Y_{i,e+l}-Y_{i,e-1}\mid E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}\mid E_{i}=\infty]$
$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e-1}(D^{e}_{i,e-1})\mid E_{i}=e]$
- $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=\infty]$
$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=e]$
- $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=\infty]$
$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e+l}(D^{\infty}_{i,e+l})\mid E_{i}=e]$
$+$ $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=e]$
- $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=\infty]$
$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e+l}(D^{\infty}_{i,e+l})\mid E_{i}=e]$
$=$ $E[(D^{e}_{i,e+l}-D^{\infty}_{i,e+l})\cdot(Y_{i,e+l}(1)-Y_{i,e+l}(0))\mid E_{i}=e]$
$=$ $CLATT_{e,e+l}\Pr(CM_{e,e+l}\mid E_{i}=e).$ (7)
The first equality follows from Assumptions 810. The second equality follows from Assumption 12. The third equality follows from Assumption 14. The final equality follows from Assumption 11.
Next, we consider the denominator of $w_{e,l}^{DID}$. From Assumptions 815, we have:
$E[D_{i,e+l}-D_{i,e-1}\mid E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}\mid E_{i}=\infty]$
$=$ $E[D^{e}_{i,e+l}-D^{e}_{i,e-1}\mid E_{i}=e]-E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=\infty]$
$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=e]-E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=\infty]$
$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e+l}\mid E_{i}=e]+E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=e]$
- $E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=\infty]$
$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e+l}\mid E_{i}=e]$
$=$ $\Pr(CM_{e,e+l}\mid E_{i}=e).$ (8)
The first equality follows from Assumption 8. The second equality follows from Assumption 12. The third equality follows from Assumption 13. The final equality follows from Assumption 11.
Combining (7) with (8), we obtain the desirable result. ∎
Proof of Theorem 3
Proof.
Fix $e\in\mathcal{E}$ and $l\in\{0,\dots,T-e\}$ such that $l<\bar{e}-e$. Note that Assumption 18 guarantees that the Wald-DID estimand $w_{e,l}^{DID,ny}$ is well defined.
First, we consider the numerator of $w_{e,l}^{DID,ny}$. From Assumptions 812 and Assumption 17, we have:
$E[Y_{i,e+l}-Y_{i,e-1}\mid E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}\mid Z_{i,e+l}=0,E_{i,e}=0]$
$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=e]$
- $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid Z_{i,e+l}=0,E_{i,e}=0]$
$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e+l}(D^{\infty}_{i,e+l})\mid E_{i}=e]$
$+$ $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid E_{i}=e]$
- $E[Y_{i,e+l}(D^{\infty}_{i,e+l})-Y_{i,e-1}(D^{\infty}_{i,e-1})\mid Z_{i,e+l}=0,E_{i,e}=0]$
$=$ $E[Y_{i,e+l}(D^{e}_{i,e+l})-Y_{i,e+l}(D^{\infty}_{i,e+l})\mid E_{i}=e]$
$=$ $E[(D^{e}_{i,e+l}-D^{\infty}_{i,e+l})\cdot(Y_{i,e+l}(1)-Y_{i,e+l}(0))\mid E_{i}=e]$
$=$ $CLATT_{e,e+l}\Pr(CM_{e,e+l}\mid E_{i}=e).$ (9)
The first equality follows from Assumptions 810 and Assumption 12. The third equality follows from Assumption 17. The final equality follows from Assumption 11.
Next, we consider the denominator of $w_{e,l}^{DID}$. From Assumptions 812 and Assumption 16, we have:
$E[D_{i,e+l}-D_{i,e-1}\mid E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}\mid Z_{i,e+l}=0,E_{i,e}=0]$
$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=e]-E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid Z_{i,e+l}=0,E_{i,e}=0]$
$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e+l}\mid E_{i}=e]+E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid E_{i}=e]$
- $E[D^{\infty}_{i,e+l}-D^{\infty}_{i,e-1}\mid Z_{i,e+l}=0,E_{i,e}=0]$
$=$ $E[D^{e}_{i,e+l}-D^{\infty}_{i,e+l}\mid E_{i}=e]$
$=$ $\Pr(CM_{e,e+l}\mid E_{i}=e).$ (10)
The first equality follows from Assumption 8 and Assumption 12. The third equality follows from Assumption 13. The final equality follows from Assumption 11.
Combining (9) with (10), we obtain the desirable result. ∎
Proof of Theorem 4
Proof.
First, we show that $\widehat{CLATT}_{e,e+l}$ is consistent for the corresponding $CLATT_{e,e+l}$. Consider case (i), where a never-exposed cohort is used as the control group. By the Law of Large Numbers (LLN) and the Continuous Mapping Theorem (CMT), $\widehat{CLATT}_{e,e+l}\xrightarrow[p]{}w^{DID}_{e,l}$ as defined in Section 3.3. Then, by Theorem 2, we have $w^{DID}_{e,l}=CLATT_{e,e+l}$. The same argument, together with Theorem 3, establishes consistency under case (ii), where not-yet-exposed cohorts are used as the control group.
@@
Next, we prove that our STS estimator is asymptotically normal, deriving its influence function. This derivation does not depend on whether we are in case (i) or case (ii). We define $\theta_{e,l}$ to be
@@
$=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}\Bigg(\frac{\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]}{E[\mathbf{1}\{E_{i}=e\}]}-\frac{\mathbf{1}\{E_{i}\in U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}\in U])\right]}{E[\mathbf{1}\{E_{i}\in U\}]}\Bigg),$ (11)
@@
$+\frac{E\big[\mathbf{1}\{E_{i}\in U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}\in U]\right]^{2}\big]}{E[\mathbf{1}\{E_{i}\in U\}]^{2}}\Bigg).$